A car of mass 1000kg moves at a consent speed of 20 m/s along the horizontal road. If  a consent frictional forces of 200N is acting between the car and the road . Calculate the power developed by engine 
 
 

Grade:11

1 Answers

Uchit Adhikari
3 years ago
Mass of car. 1000kg
Acceleration of car 0
Resultant force =m*a
                                1000*0=0
Resultant force(Fr)=Fe-Ff
                       0=Fe - 200N
Fe=200N